The volume of a rectangular box is 252 cu. units. The height is 4 units and the length is two more than the width of the box. What are the width and length of the box?
Thank you so much for your assistance

Answer by checkley77(12844)

H=4
L=W+2
V=LWH
252=(W+2)W*4
252=(W^2+2W)4
252=4W^2+8W
4W^2+8W-252=0
4(W^2+2W-63)=0
4(W+9)(W-7)=0
W-7=0
W=7 ANS. FOR THE WIDTH.
L=7+2=9 ANS. FOR THE LENGTH.
PROOF:
252=9*7*4
252=252